Define elasticity of supply and find the price from the given statement:

If Es of a good is 2 and a firm supplies 200 units at price of Rs 8 per unit, then at what price will the firm supply 250 units. What is the conclusion?

1
Expert's answer
2021-05-17T10:32:20-0400

We define elasticity of supply of a good as the amount of the change in the quantity demanded of a good due to one percent change in the good's price.


Elasticity of supply "=\frac{\delta Q}{\delta P}=\frac{\frac{Q_2-Q_1}{Q1}}{\frac{P_2-P_1}{P_1}}"


Given : Elasticity of supply (Es)"= 2 \space \space"

"0.125 = \frac{(X - 8)}{8}"

P1 = 8  and Q1 = 200

P2 = ? and Q2 = 250

% change in quantity supplied"= \frac{(250−200)}{200} = 0.25 = 25\%"


so, using formula , we have :"2 = \frac{25\%}{\% \space \delta\space in P}"

% change in P"=\frac{ (P2 − 8) ×100}{8}= \frac{25\%}{2} = 12.5\%"


"(P_2 \u2212 8)\u00d7100 = 100"

"(P_2 − 8) =\frac{ 100}{100} = 1"

"P_2 \u2212 8 = 1"

"P_2 = 1 + 8 = 9"


Thus, "P_2" is Rs 9 or the price is Rs 9 when the quantity sold is 250 units.

 

The conclusion is that the supply of the good is elastic as the Es is more than 1 and any change in price causes more than proportional change in quantity demanded.